OpenStudy (swamijatin2):

find the value of n 9^n+9^n+9^n = 3^2013

OpenStudy (swamijatin2):

need help

OpenStudy (anonymous):

9^n+9^n+9^n = 3^2013 3*9^n = 3^2013 9^n = 3^2012 3^(2n) = 3^(2012) 2n = 2012 n=1006

OpenStudy (swamijatin2):

i had problem with second step

OpenStudy (anonymous):

?

OpenStudy (swamijatin2):

9^n+9^n+9^n = 3^2013 3*9^n = 3^2013 how this is done

OpenStudy (anonymous):

let's call 9^n : x so you have x + x + x = 3^2013 3x = 3^2013 now what is x ? 3*9^n = 3^2013

OpenStudy (swamijatin2):

oh i got it thanxs ............